A university claims that the percentage of its graduates who fail to get a job after one year of graduation is less than 10%. A sample of 200 graduates from the university was randomly selected and found that 16 graduates fail to get a job after one year of graduation. Is there enough evidence to support the university's claim? Test at 5% significance level. 


1
Expert's answer
2022-01-17T16:27:18-0500

Sample Proportion "\hat{p}=\dfrac{16}{200}=0.08"

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\geq0.1"

"H_1:p<0.1"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a left-tailed test is "z_c = -1.6449."

The rejection region for this two-tailed test is "R = \{z: z<-1.6449\}."

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.08-0.1}{\sqrt{\dfrac{0.1(1-0.1)}{200}}}"

"\approx-0.9428"

Since it is observed that "z=-0.9428> -1.6449=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=P(Z<-0.9428)=0.1728916," and since "p=0.1728916>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population proportion "p" is less than "0.1," at the "\alpha = 0.05" significance level.

Therefore, there is not enough evidence to claim that the percentage of the graduates who fail to get a job after one year of graduation is less than 10%, at the "\alpha = 0.05" significance level.